What is the value of x in the equation 3x-4y=65 when y=4


1
Expert's answer
2021-09-30T15:39:31-0400

given

"3x-4y=65"

and "y=4"

we substitute for y in the equation

"3x-4(4)=65\\3x=65+16\\3x=81\\x=27"